# Local Identification in Instrumental Variable Multivariate Quantile Regression Models

### v3

@@ 2.3 Comparison with [chernozhukov2005iv]
More importantly, the dimension-wise approach typically leads to the violation of the rank similarity assumption of [chernozhukov2005iv]. For consider a binary treatment environment and assume that satisfies Assumption 2.1 with the rank invariance and where is a random variable independent of all the other variables. The potential outcome for treatment is componentwisely represented as
@@
Since and are independent, it is expected that the treatment effect of is identified if is sufficiently informative.
@@
Consider an empirical researcher who is interested only in the first component of the potential outcome. In the Engel curve example, this describes the situation where the researcher is interested only in the expenditure share of food. In this case, it is common to apply the standard quantile model of [chernozhukov2005iv] to the data To see that this practice leads to the violation of the rank similarity assumption, notice that the structural equation under the one-dimensional model is where and is the quantile function of and that is measurable in via The rank similarity requires but this is satisfied only when which does not hold in general.222For a counterexample, consider and In other words, the instrument is invalid in the sense that it is correlated with the unobserved heterogeneity via the treatment variable
The failure of the rank similarity is caused by the fact that depends not only on but also on where and [chernozhukov2005iv] do not allow such dependence, which is often unrealistic. Specifically, in order to apply [chernozhukov2005iv] to the Engel curve dataset in the previous subsection, it must be assumed that individuals choose the expenditure share on food based only on their preference for food. However, they should also care about their preference for leisure, since it determines how much they can spend on food. Our multivariate quantile model allows for the dependence of the potential outcomes on the multiple unobservables which remedies the problem.
@@
2.4 Connections to optimal transport theory
Our multivariate outcome model is tightly related to the optimal transport theory. By Assumption 2.1, the structural function transports to which is denoted by where is the distribution of that is, it satisfies for all measurable Also, remember that has a symmetric and positive definite derivative. It is well-known that is optimal in the sense that it minimizes the quadratic transportation cost; that is, solves
where is the Euclidean norm, and it is called the optimal transport map from to
It is also known that the optimal transport map characterizes the target distribution once the source distribution is fixed. It also has a certain monotonicity property, called the cyclical monotonicity. Recent developments in statistical theory have shown that the optimal transport map can be thought of as a multivariate extension of quantile functions (see, e.g., [ekeland2012comonotonic], [chernozhukov2017monge], [hallin2021distribution], [ghosal2022multivariate]). In this literature, the optimal transport map from to is also called the multivariate quantile function of with respect to the reference measure Following this terminology, we call our model an IV multivariate quantile regression model.
See Appendix A for more details.
@@ 3 Identification
The representation (1) of Theorem 2.1 and the change-of-variables formula imply that the true structural functions solve the functional equations
@@
The conditions in the definition in restricts possible deviations from the true structural function. This type of condition is called a source condition and usually assumed in the literature of inverse problems and nonparametric IV identification. See, for example, [carrasco2007linear], [chen2014local], [scaillet2016ill], and [centorrino2024iterative]. Section 2.3 of [chen2014local] discusses an example in which identification fails if an analogous condition is dropped.
@@
Assumption 3.2 requires that the support of a potential outcome be regular. The convexity of the support is necessary for many purposes in the optimal transport theory (see, for example, [villani2003topics] and [figalli2017monge]). The smoothness of the boundary is rarely a problem in practice, as it allows for kinks at some points. Combined with Assumption 3.2, Assumption 3.3 implies that is bounded.
@@
Condition (4) is a novel assumption that we cannot find in the literature. Roughly speaking, this condition requires () to be sufficiently positively correlated with (). Too see this, let be the density of conditional on and and suppose that it satisfies for some constants Then, the condition
is sufficient for condition (4). This holds if and are large enough. Since condition (4) requires a positive correlation between and the labeling of the IV should be switched in favor of the condition.
4 Discussions
4.1 Cost of multidimensionality
@@
Condition (4) is stronger than matrix (5) being positive definite, which can be shown as follows. Since equation (4) implies
and therefore, the determinant of matrix (5) is strictly positive. As its trace is clearly positive, it is positive definite.
Proposition 4.1 is consistent with [chernozhukov2005iv], who show that the identification holds if matrix (5) is full-rank, because the determinant of any full-rank matrix can be positive by switching columns if necessary.
The identification for holds under a milder assumption because the proof of Proposition 4.1 uses a fact that is valid only in dimension one. Specifically, it depends on the fact that the cofactor matrix333For an invertible matrix its cofactor matrix is defined as of any matrix is the identity if only if See Lemma C.4 for details. This implies that the positive definiteness of matrix (5) being replaced by condition (4) can be seen as the cost of taking multidimensional potential outcomes into account.
As the dimension increases, condition (4) becomes more demanding. This is quite natural, as one needs to identify higher dimensional structural functions. However, the minimum size of the support of the IV, needed for the identification is independent of the dimension of the potential outcomes, as long as the IV is informative enough.
@@
Roughly speaking, the proof of Theorem 3.1 consists of two steps. In the first step, we show that the support of the potential outcome is identified under condition (4). If a candidate of the structural function induces a different support of the potential outcome, then it cannot be consistent with the joint distribution of the observable variables Thus, we can focus on the identification in the set of structural functions that preserve the support.
In the second step, we show that is the locally unique solution to the system (3) of equations in the set of structural functions. We use an implicit-function-theorem-type argument that is common in the literature on the identification of finite-dimensional parameters in nonlinear models. We consider the linearization of the nonlinear system (3) around the truth It turns out that the positive correlation condition (4) implies that the “slope” of the linearized system is away from zero, which shows the local uniqueness of the solution.
@@ Appendix A Preliminaries
In this section, we briefly review mathematical concepts that are used in the paper. Fix a probability measure on a compact convex subset of the -dimensional Euclidean space, and let be a -dimensional random vector. First, consider the case of and According to the inverse probability integral transform, there exists a random variable such that almost surely where is the quantile function of (see, for example, Proposition 3.2 of [shorack2000probability]). Moreover, is the (almost surely) unique nondecreasing function that satisfies this equation for some uniform random variable. This observation for the classical one-dimensional case holds more generally by replacing the monotonicity of with the restriction that it is written as the gradient of a convex function.
@@
This is a version of the fundamental theorem in the optimal transport theory known as the Brenier-McCann theorem ([brenier1991polar] and [mccann1995existence]). The gradient is called the optimal transport map from to the distribution of under the quadratic cost because it solves the following Monge problem:
@@
as is shown in [rockafellar1966characterization]. Conversely, if a function has a symmetric and positive definite derivative, then it is written as the gradient of a convex function.
@@
Furthermore, a multivariate quantile function induces a bijection between the interior of the domain and that of the range under some regularity conditions. For similar results, see Theorem 1 of [cordero2019regularity] and Proposition 3.1 of [ghosal2022multivariate].
@@ B.1 Setup and theorem
This result is closely related to Section 2 of [chen2014local]. Indeed, their Theorems 1 can be shown from Theorem B.1. Our theorem has a different focus than their result in that it covers the local identification problem for systems of equations and that the domain of functionals is not assumed to be a Banach space.
@@
We first show an equivalent representation of Assumption B.2.
@@ Appendix E Omitted proofs
Proof.
The optimization problem (2) can be written as
where the new variable can be understood as the total expenditure. The inner maximization problem is written as
where and Note that is the budget share of good By eliminating this problem is equivalent to
where and which is concave in Here, we can think of as the consumer’s relative propensity to consume good Under the assumption that this problem has an interior solution, if the consumer has consumption propensity and the total expenditure is the optimal shares of goods are given by
where is the convex conjugate of ([rockafellar1970convex]). Since is convex, the argument above implies that the vector of the optimal expenditure shares has a symmetric and positive semidefinite derivative, if it exists. ∎
E.3 Proof of Proposition 4.1
E.4 Proof of Theorem A.1
@@
As is symmetric on which is a convex domain, there exists such that by Poincaré’s lemma. Since is positive definite, is strictly convex. The uniqueness follows immediately from Lemma 2.1 of [del2019central]. ∎
E.6 Proof of Theorem A.2
@@
where the last equality follows from Piola’s identity. See, for example, the lemma in page 440 of [evans2010partial]. ∎
@@
The proof of this lemma is inspired by [delanoe1991classical] and [urbas1997second]. Let be a point. There exists a neighborhood of such that is Then, by Assumption 3.2 and a version of Theorem 5.6 of [delfour1994shape], there exists a continuously differentiable function such that and for Notice that is an outward normal vector of at Let Also, there is such that To see this, recall that as is compact. Thus, there is such that As we have by Lemma E.3.

### v4

@@ 2.3 Comparison with [12]
Even when multiple outcome variables are present, one might consider applying the one-dimensional quantile model of [12] to each component separately. However, this approach typically leads to the violation of the rank similarity assumption of [12]. To see this, we revisit the wealth accumulation example of Section 2.2. For consider a binary treatment environment and assume that satisfies Assumption 2.1 with the rank invariance and where is a random variable independent of all the other variables. Recall that is the vector consisting of net financial assets and other assets, is the participation status of 401(k), is its eligibility, and is the preference for the corresponding assets. The potential outcome for treatment is componentwisely represented as
@@
Since and are assumed independent, it is expected that the treatment effect of is identified if is sufficiently informative.
@@
Consider an empirical researcher who is interested only in net financial assets In this case, it is common to apply the standard quantile model to the data as [11] do. To see that this practice leads to the violation of the rank similarity assumption, notice that the structural equation under the one-dimensional model is where and is the quantile function of Then, is measurable in since The rank similarity requires but this is satisfied only when which does not hold in general.222For a counterexample, consider and In other words, the instrument is invalid in the sense that it is correlated with the unobserved heterogeneity via the treatment variable
The failure of rank similarity arises because depends not only on but also on whereas [12] do not allow such dependence—an assumption that is often unrealistic. Specifically, applying the framework of [12] to the wealth accumulation example would require assuming that individuals determine their holdings of financial assets solely based on their preference for financial assets. In reality, individuals also consider their preference for other assets, since it affects how much they can allocate to financial assets. Our multivariate quantile model accommodates such interdependence by allowing the potential outcomes to depend jointly on multiple unobservables thereby resolving this limitation.
@@ 3 Identification
The representation (1) of Theorem 2.1 and the change-of-variables formula imply that the true structural function solves the measure-valued equation
@@
The conditions in the definition of restrict the class of admissible deviations from the true structural function. Such restrictions are common in the literature on nonparametric IV identification. To see their role, note that identifying requires inverting the linearization of the system of equations (2). Although this linearized system is invertible under the standard full-rank condition, its inverse is generally discontinuous because of the infinite-dimensional nature of the problem, rendering the identification ill-posed. This difficulty is resolved by restricting the identification domain to thereby ruling out pathological deviations from the truth. Similar regularity conditions are imposed in related work, including [8] and [7]. In particular, Section 2.3 of [8] presents an example where identification fails once an analogous restriction is removed.
@@
Assumption 3.2 requires that the support of a potential outcome be regular. The convexity of the support is necessary for many purposes in the optimal transport theory (see, for example, [39] and [23]). The smoothness of the boundary is rarely a problem in practice, as it allows for kinks at some points. Combined with Assumption 3.2, Assumption 3.3 implies that is bounded.
@@
Condition (3) is a novel assumption that, to the best of our knowledge, has not appeared in the existing literature. Roughly speaking, it requires that () be sufficiently positively correlated with (). To see this, let denote the density of conditional on and and suppose it satisfies for some constants Then, the condition
is sufficient for condition (3) to hold. This inequality is satisfied when and are sufficiently large. Since condition (3) requires a positive correlation between and the labeling of the instrumental variable can be reversed if necessary to meet the condition.
Condition (3) is stronger than [12]’s monotone likelihood ratio condition:
This can be shown as follows. Since equation (3) implies
@@
Proposition 3.1 is consistent with [12], who show that identification holds when matrix (4) is full rank, since the determinant of any full-rank matrix can be made positive by switching its columns if necessary.
Identification for holds under a milder assumption because the proof of Proposition 3.1 relies on a property that is valid only in one dimension. Specifically, it uses the fact that the cofactor matrix333For an invertible matrix , its cofactor matrix is defined as . of any matrix equals the identity if and only if . See Lemma C.4 for details. This observation implies that replacing the positive definiteness of matrix (4) with condition (3) can be interpreted as the additional cost of allowing for multidimensional potential outcomes.
As the dimension increases, condition (3) becomes more demanding, which is natural given the need to identify higher-dimensional structural functions. However, the minimum support size of the instrumental variable required for identification remains independent of the dimension of the potential outcomes, as long as the IV is sufficiently correlated with the treatment variable.
@@
Roughly speaking, the proof of Theorem 3.1 proceeds in two steps. In the first step, we show that the support of the potential outcome is identified under condition (3). If a candidate structural function implies a different support for the potential outcome, then it cannot be consistent with the joint distribution of the observable variables Therefore, the identification analysis can be restricted to the class of structural functions that preserve the support.
In the second step, we show that is the locally unique solution to the system (2) of equations within the class of structural functions. The argument follows the logic of the implicit function theorem, which is commonly used in the identification analysis of nonlinear models with finite-dimensional parameters. Specifically, we linearize the nonlinear system (2) around the true structural function The key insight is that the positive correlation condition (3) ensures that the “slope” of the linearized system is bounded away from zero, thereby guaranteeing the local uniqueness of the solution.
4 Concluding Remarks
In this paper, we proposed a new nonlinear IV model that extends [12] to accommodate the correlation among multiple outcome variables. A key identifying restriction is that the structural functions are assumed to have a symmetric and positive definite derivative with respect to the rank vector. We showed that if the instrumental variable is sufficiently positively correlated with the treatment variable, the structural functions are locally identified. The cost of multidimensionality is that the positive correlation condition becomes more demanding as the dimension of the potential outcomes increases. Nevertheless, the minimum support size of the instrumental variable required for identification remains the same as in the univariate case. These results clarify how identification in nonlinear IV models can be extended to multidimensional settings while maintaining a similar structure to the scalar case.
@@ Appendix A Preliminaries
Our multivariate outcome model is tightly related to the optimal transport theory. In this section, we briefly review mathematical concepts that are used in the paper. Fix a probability measure on a compact convex subset of the -dimensional Euclidean space, and let be a -dimensional random vector. First, consider the case of and According to the inverse probability integral transform, there exists a random variable such that almost surely where is the quantile function of (see, for example, Proposition 3.2 of [36]). Moreover, is the (almost surely) unique nondecreasing function that satisfies this equation for some uniform random variable. This observation for the classical one-dimensional case holds more generally by replacing the monotonicity of with the restriction that it is written as the gradient of a convex function.
@@
This is a version of the fundamental theorem in the optimal transport theory known as the Brenier-McCann theorem ([4] and [32]). The gradient is called the optimal transport map from to the distribution of under the quadratic cost because it solves the following Monge problem:
@@
as is shown in [34]. Conversely, if a function has a symmetric and positive definite derivative, then it is written as the gradient of a convex function.
@@
Furthermore, a multivariate quantile function induces a bijection between the interior of the domain and that of the range under some regularity conditions. For similar results, see Theorem 1 of [17] and Proposition 3.1 of [26].
@@ B.1 Setup and theorem
This result is closely related to Section 2 of [8]. Indeed, their Theorems 1 can be shown from Theorem B.1. Our theorem has a different focus than their result in that it covers the local identification problem for systems of equations and that the domain of functionals is not assumed to be a Banach space.
@@
We use the following lemma that gives an equivalent representation of Assumption B.2.
@@ Appendix E Omitted proofs
E.3 Proof of Theorem A.1
@@
As is symmetric on which is a convex domain, there exists such that by Poincaré’s lemma. Since is positive definite, is strictly convex. The uniqueness follows immediately from Lemma 2.1 of [18]. ∎
E.5 Proof of Theorem A.2
@@
where the last equality follows from Piola’s identity. See, for example, the lemma in page 440 of [22]. ∎
@@
The proof of this lemma is inspired by [19] and [38]. Let be a point. There exists a neighborhood of such that is Then, by Assumption 3.2 and a version of Theorem 5.6 of [20], there exists a continuously differentiable function such that and for Notice that is an outward normal vector of at Let Also, there is such that To see this, recall that as is compact. Thus, there is such that As we have by Lemma E.3.